OpenStudy (anonymous):

can someone help me with this worksheet will give medal for who ever gets it right @Quan99 @mathmath333 @Mr.Man420 @Wolverine32 @sarahprincess98 @PlanSlam @youngnephew34 @YOUNGSUPERMAN @Undeadwarriorprincess @Hero

OpenStudy (anonymous):

hello

OpenStudy (anonymous):

heres the problem

OpenStudy (youngsuperman):

wat u need

OpenStudy (anonymous):

sorry not good at math

OpenStudy (anonymous):

help ! @YOUNGSUPERMAN

OpenStudy (youngsuperman):

yes

OpenStudy (anonymous):

look at the problem

OpenStudy (anonymous):

@YOUNGSUPERMAN

OpenStudy (youngsuperman):

kk

OpenStudy (anonymous):

do you think you can solve it ... :/ @YOUNGSUPERMAN

OpenStudy (youngsuperman):

ion know it hard

OpenStudy (anonymous):

-_- ugh @YOUNGSUPERMAN

OpenStudy (youngsuperman):

im sorry pm me

OpenStudy (anonymous):

@iGreen. can you herlp me

OpenStudy (mathmath333):

\(\LARGE \begin{align} \color{black}{ \sqrt[6]{125a^{15}} \hspace{.33em}\\~\\ =\left(125a^{15}\right)^{\frac{1}{6}} \hspace{.33em}\\~\\ =\left(5^3a^{15}\right)^{\frac{1}{6}} \hspace{.33em}\\~\\ =5^{\frac{3}{6}}a^{\frac{15}{6}} \hspace{.33em}\\~\\ =5^{\frac{1}{2}}a^{\frac{5}{2}} \hspace{.33em}\\~\\ }\end{align}\)

OpenStudy (anonymous):

@nobody_need_know97

OpenStudy (mathmath333):

\(\huge y = a(x – h)^2 + k\)

OpenStudy (mathmath333):

what is h and k compared to ur equation

OpenStudy (mathmath333):

here a=2

OpenStudy (mathmath333):

can h=0 and k=0 ?

OpenStudy (anonymous):

um yes @mathmath333

OpenStudy (mathmath333):

yes so h=0 and k=0

OpenStudy (mathmath333):

vertex\((0,0)\)

OpenStudy (anonymous):

ok heres some more

OpenStudy (anonymous):

@mathmath333

OpenStudy (anonymous):

i think i know that

OpenStudy (mathmath333):

formula for focus is this \(\large (h, k + \dfrac{1}{(4a)})\)